# On the Convergence Analysis of Muon

### v1

@@ 4.1 Nonconvex case
Denote $\Delta=f(W_{0})-f^{*}$. If we set $B=1$, $\eta=\sqrt{\frac{(1-\beta)\Delta}{rTL}}$, $1-\beta=\min\{\frac{\sqrt{L\Delta}}{\sigma\sqrt{T}},1\}$, then
@@
Thus, we can find an $\epsilon$-nuclear norm stationary point of $f$ with a complexity of $O(r^{2}L\sigma^{2}\Delta\epsilon^{-4})$.
Note that if we set $\beta=O(1)$ (or $\beta=0$) in the stochastic setting, a large batch size $B=O(\epsilon^{-2})$ is required to sufficiently reduce the variance of the stochastic mini-batch gradient. Therefore, our theoretical analysis suggests that momentum can be beneficial in the stochastic setting, which aligns with the observation in Figure 1 where Muon with momentum outperforms Muon without momentum in stochastic setting.
We have the following corollary for the deterministic setting.
Corollary 4.2.
Under Assumption 3.1, if we apply Algorithm 1 with $G_{t}=\nabla f(W_{t})$ and $\eta_{t}=\eta$, then
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla f(W_{t})\|_{*}\leq\frac{f(W_{0})-f(W_{T})}{T\eta}+\frac{Lr\eta}{2}+\frac{2r\eta\beta L}{1-\beta}$
Denote $\Delta=f(W_{0})-f^{*}$. If we set $\eta=\sqrt{\frac{\Delta}{rTL}}$, $\beta=O(1)$, then
@@
Thus, we can find an $\epsilon$-nuclear norm stationary point of $f$ with a complexity of $O(rL\Delta\epsilon^{-2})$.
Our theoretical results indicate that in the deterministic setting, setting $\beta=O(1)$ (or even $\beta=0$) does not significantly affect the convergence guarantee. This is consistent with the empirical observation in Figure 1, where Muon with momentum performs similarly to Muon without momentum in the deterministic setting.
Note that the (S)GD’s complexity of finding $\epsilon$-Frobenius norm stationary points of $f$ is $O(L\sigma^{2}\Delta\epsilon^{-4})$ in the stochastic setting and $O(L\Delta\epsilon^{-2})$ in the deterministic setting (Garrigos and Gower, 2023), which means (S)GD can also find $\epsilon$-nuclear norm stationary points of $f$ with $O(r^{2}L\sigma^{2}\Delta\epsilon^{-4})$ complexity in the stochastic setting and $O(rL\Delta\epsilon^{-2})$ in the deterministic setting according to Proposition 3.6. Thus, there is little direct advantage of Muon over (S)GD when analyzing the convergence rate with the conventional Frobenius norm Lipschitz smoothness assumption.
However, if we adopt the spectral norm Lipschitz smoothness assumption, the comparison changes.
@@
Denote $\Delta=f(W_{0})-f^{*}$. If we set $B=1$, $\eta=\sqrt{\frac{(1-\beta)\Delta}{TL_{*}}}$, $1-\beta=\min\{\frac{\sqrt{L_{*}\Delta}}{\sigma\sqrt{rT}},1\}$, then
@@
Thus, we can find an $\epsilon$-nuclear norm stationary point of $f$ with a complexity of $O(rL_{*}\sigma^{2}\Delta\epsilon^{-4})$.
Similarly, we have the following corollary for the deterministic setting.
Corollary 4.4.
Under Assumption 3.2, if we apply Algorithm 1 with $G_{t}=\nabla f(W_{t})$ and $\eta_{t}=\eta$, then
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla f(W_{t})\|_{*}\leq\frac{f(W_{0})-f(W_{T})}{T\eta}+\frac{L_{*}\eta}{2}+\frac{2\eta\beta L_{*}}{1-\beta}.$
Denote $\Delta=f(W_{0})-f^{*}$. If we set $\eta=\sqrt{\frac{\Delta}{TL_{*}}}$, $\beta=O(1)$, then
@@
Thus, we can find an $\epsilon$-nuclear norm stationary point of $f$ with a complexity of $O(L_{*}\Delta\epsilon^{-2})$.
For a function $f$, its $L$ and $L_{*}$ are related to its Hessians $\nabla^{2}f$. In neural network optimization, a widely observed phenomenon is that Hessians are typically low-rank (Sagun et al., 2016, 2017; Wu et al., 2020; Papyan, 2020; Yao et al., 2020) and are approximately blockwise diagonal (Dong et al., 2025; Zhang et al., 2024b, a; Collobert, 2004) (e.g., See Figure 2 in Zhang et al. (2024a) and Figure 1-3 in Dong et al. (2025)). Thus, we make the following assumption to better illustrate the connections and differences between $L$ and $L_{*}$ in neural networks.
@@
We say $f:\mathbb{R}^{m\times n}\to\mathbb{R}$ is 1-$\Lambda$-norm Lipschitz smooth if for any $W,W^{\prime}\in\mathbb{R}^{m\times n}$,
@@
where $\Lambda\in\mathbb{R}^{n\times n}$ is a positive definite matrix.
4.5 is closely related to the blockwise diagonal structure of the Hessian and is equivalent to the the following assumption of Hessians:
@@
where $f_{v}(w)=f(W)$ and $w=\mathsf{vec}(W)\in\mathbb{R}^{mn}$. 4.5 has also been used in An et al. (2025) for analyzing matrix-structured optimization.
@@
Thus, if the Hessians of $f$ have a blockwise diagonal structure and are relatively “low rank”, i.e. if it satisfies 4.5 with $\Lambda$ such that $\|\Lambda\|_{{\rm op}}\approx\|\Lambda\|_{*}\ll r\|\Lambda\|_{{\rm op}}$, then the convergence rate of Muon for finding the nuclear norm stationary points are better than (S)GD’s, i.e. $O(\|\Lambda\|_{*}\Delta\epsilon^{-2})\ll O(r\|\Lambda\|_{{\rm op}}\Delta\epsilon^{-2})$ and $O(r\|\Lambda\|_{*}\sigma^{2}\Delta\epsilon^{-4})\ll O(r^{2}\|\Lambda\|_{{\rm op}}\sigma^{2}\Delta\epsilon^{-4})$ for deterministic and stochastic settings respectively.
@@
To further investigate how the structure and dynamics of the Hessian matrix affect the convergence of Muon and GD in a more fine-grained manner, we consider the following assumption and analysis.
@@
We assume the norm of $\nabla^{3}f$ is bounded, i.e., we assume for any $W\in\mathbb{R}^{m\times n}$, $\|\nabla^{3}f_{v}(w)\|_{{\rm op}}\leq s$. Here $f_{v}(w)=f(W)$ and $w=\mathsf{vec}(W)\in\mathbb{R}^{mn}$.
As discussed before, the momentum term is mainly used to reduce the variance of the stochastic mini-batch gradient. It plays a similar role as the standard SGD with momentum. The key difference between Muon and GD is the modified update direction – the orthogonal update direction. Thus, for simplicity, we consider Algorithm 2, the Muon without momentum in the deterministic setting.
@@
Under 4.7, if we apply Algorithm 2 with $\eta_{t}=\eta$, then
@@
[…]
Thus, we can find an $\epsilon$-nuclear norm stationary point of $f$ with a complexity of $O(J\Delta\epsilon^{-2})$.
Note that
$J_{t}$ $=\mathsf{vec}(U_{t}V_{t}^{\top})^{\top}H_{t}\mathsf{vec}(U_{t}V_{t}^{\top})$
$=\mathsf{vec}(I_{r_{t}})^{\top}(U_{t}^{\top}\otimes V_{t}^{\top})H_{t}(U_{t}\otimes V_{t})\mathsf{vec}(I_{r_{t}})$

### v2

@@ 4.1 Frobenius norm Lipschitz smooth
By choosing the parameters in Theorem 4.1, we have following corollary.
Corollary 4.2.
Under the same assumptions as Theorem 4.1, we have the following convergence guarantees for Algorithm 1:
  1. 1.
    Stochastic setting: If we set $B=1$, $\eta=\sqrt{\frac{(1-\beta)\Delta}{rTL}}$, and $1-\beta=\min\left\{\frac{\sqrt{L\Delta}}{\sigma\sqrt{T}},1\right\}$, then
@@
  2. 2.
    Deterministic setting ($\sigma=0$): If we set $\eta=\sqrt{\frac{\Delta}{rTL}}$, and $\beta=O(1)$, then
@@
Therefore, the complexity to find an $\epsilon$-nuclear norm stationary point of $f$ is $O(r^{2}L\sigma^{2}\Delta\epsilon^{-4})$ in the stochastic setting and $O(rL\Delta\epsilon^{-2})$ in the deterministic setting.
Moreover, to study the convergence of function value of Muon, we also consider the star convex condition.
Assumption 4.3 (Star convex).
$\langle\nabla f(W),W-W^{*}\rangle\geq f(W)-f^{*}.$
where $W^{*}\in\mathop{\rm argmin}_{W}f(W)$ is the optimal point.
To ensure the convergence of Muon in the star convex case, we adopt the following mild assumption. Note that the same assumption is also used in Gupta et al. (2018); An et al. (2025) for their convergence analysis of their algorithms in similar settings.
Assumption 4.4.
For $W_{t},t=0,\dots,T$, generated by Algorithm 2, we assume $\|W_{t}-W^{*}\|_{{\rm op}}\leq D_{{\rm op}}$.
In the experiment (Figure 1(b)), we can observe that when we use Muon on a quadratic function, $\|W_{t}-W^{*}\|$ is bounded, which satisfies 4.4.
With these assumptions, we have the following theorem.
Theorem 4.5.
Under Assumptions 4.3, 3.1, and 4.4, if we apply Algorithm 2 with $\eta_{t}=\eta$, then
Thus, to reach the precision $f(W_{T})-f^{*}\leq\epsilon$, we can take $\eta=O(\frac{\epsilon}{rLD_{{\rm op}}})$, and the complexity is $O(rLD_{{\rm op}}^{2}\epsilon^{-1}\text{log}\frac{\Delta}{\epsilon})$.
Theorem 4.6.
Under the same assumptions as Theorem 4.5, if we apply Algorithm 2 with adaptive learning rates $\eta_{t}=\frac{\|\nabla f(W_{t})\|_{*}}{rL}$, we have
4.1.1 Comparison with GD
Note that in the nonconvex case, the (S)GD’s complexity of finding $\epsilon$-Frobenius norm stationary points of $f$ is $O(L\sigma^{2}\Delta\epsilon^{-4})$ in the stochastic setting and $O(L\Delta\epsilon^{-2})$ in the deterministic setting (Garrigos and Gower, 2023), which means (S)GD can also find $\epsilon$-nuclear norm stationary points of $f$ with $O(r^{2}L\sigma^{2}\Delta\epsilon^{-4})$ complexity in the stochastic setting and $O(rL\Delta\epsilon^{-2})$ in the deterministic setting according to Proposition 3.6. Moreover, to reach the precision $f(W_{T})-f^{*}\leq\epsilon$ in the star convex case, the complexity of GD is $O(LD_{F}^{2}\epsilon^{-1})$ (Garrigos and Gower, 2023). We have $D_{F}=\|W_{0}-W^{*}\|_{F}\leq\|W_{0}-W^{*}\|_{*}\leq D_{\rm op}$. Comparing the complexity results in Theorem 4.1 and Theorem 4.5, we can observe that under the conventional Frobenius norm Lipschitz smoothness assumption, the convergence rate of Muon is no better than that of (S)GD, and may even be worse.
4.2 Spectral norm Lipschitz smooth
Considering the matrix structure of the parameters and the properties of the Muon optimizer, the spectral norm Lipschitz smoothness assumption is natural in practice. In this subsection, we adopt 3.2 and establish the following theorems and corollaries.
@@ 4.2 Spectral norm Lipschitz smooth
Corollary 4.8.
Under the same assumptions as Theorem 4.7, we have the following convergence guarantees for Algorithm 1:
  1. 1.
    Stochastic setting: If we set $B=1$, $\eta=\sqrt{\frac{(1-\beta)\Delta}{TL_{*}}}$, and $1-\beta=\min\left\{\frac{\sqrt{L_{*}\Delta}}{\sigma\sqrt{rT}},1\right\}$, then
@@
  2. 2.
    Deterministic setting ($\sigma=0$): If we set $\eta=\sqrt{\frac{\Delta}{TL_{*}}}$, and $\beta=O(1)$, then
@@
Consequently, achieving an $\epsilon$-nuclear norm stationary point of $f$ requires a complexity of $O(rL_{*}\sigma^{2}\Delta\epsilon^{-4})$in stochastic setting and $O(L_{*}\Delta\epsilon^{-2})$ in deterministic setting.
Theorem 4.9.
Under 4.3, 3.2 and 4.4, if we apply Algorithm 2 with $\eta_{t}=\eta$, then
Thus, to reach the precision $f(W_{T})-f^{*}\leq\epsilon$, we can take $\eta=O(\frac{\epsilon}{L_{*}D_{{\rm op}}})$, and the complexity is $O(L_{*}D_{{\rm op}}^{2}\epsilon^{-1}\text{log}\frac{\Delta}{\epsilon})$.
Theorem 4.10.
Under the same assumptions as Theorem 4.9, if we apply Algorithm 2 with adaptive learning rates $\eta_{t}=\frac{\|\nabla f(W_{t})\|_{*}}{L_{*}}$, we have
Though we have derived the convergence theorems in terms of $L_{*}$, it remains unclear how to interpret them. For the Frobenius norm Lipschitz smoothness constant $L$, we know it is the largest singular value of the Hessian. The question, therefore, is what the relationship between $L$ and $L_{*}$ is. For a function $f$, its $L$ and $L_{*}$ are related to its Hessian $\nabla^{2}f$. In neural network optimization, a widely observed phenomenon is that Hessians are typically low-rank (Sagun et al., 2016, 2017; Wu et al., 2020; Papyan, 2020; Yao et al., 2020) and are approximately blockwise diagonal (Dong et al., 2025; Zhang et al., 2024b, a; Collobert, 2004) (e.g., See Figure 2 in Zhang et al. (2024a) and Figure 1-3 in Dong et al. (2025)). Thus, we adopt 4.11 from An et al. (2025) to better illustrate the connections and differences between $L$ and $L_{*}$ in neural networks.
@@
We say $f:\mathbb{R}^{m\times n}\to\mathbb{R}$ is 1-$\Lambda$-norm Lipschitz smooth with a positive definite matrix $\Lambda\in\mathbb{R}^{n\times n}$, if for any $W,W^{\prime}\in\mathbb{R}^{m\times n}$,
@@
4.11 is closely related to the blockwise diagonal structure of the Hessian. Actually, when $f$ is twice continuously differentiable, 4.11 is equivalent to the following assumption of Hessian:
@@
where $f_{v}(w)=f(W)$ and $w=\overline{\mathrm{vec}}(W)\in\mathbb{R}^{mn}$. The proof of the equivalence of these two assumptions can be found in Lemma A.3. 4.11 has also been used in An et al. (2025) to analyze matrix-structured optimization. Moreover, we have the following lemma to establish the relationship between Assumptions 3.1, 3.2, and 4.11, whose proof can be found in Appendix A.
@@ 4.2.1 Comparison with GD
In this section, we first argue that Muon inherently exploits blockwise diagonal structures and relatively low-rank properties. For example, if the Hessians of $f$ have a blockwise diagonal structure and are relatively “low rank”, i.e. if it satisfies 4.11 with $\Lambda$ such that $\|\Lambda\|_{{\rm op}}\approx\|\Lambda\|_{*}\ll r\|\Lambda\|_{{\rm op}}$, then the convergence rate of Muon for finding the nuclear norm stationary points can be better than that of (S)GD in the nonconvex case, i.e. $O(\|\Lambda\|_{*}\Delta\epsilon^{-2})\ll O(r\|\Lambda\|_{{\rm op}}\Delta\epsilon^{-2})$ and $O(r\|\Lambda\|_{*}\sigma^{2}\Delta\epsilon^{-4})\ll O(r^{2}\|\Lambda\|_{{\rm op}}\sigma^{2}\Delta\epsilon^{-4})$.
In the star convex case, recall that the complexity of GD is $O(LD_{\rm F}^{2}\epsilon^{-1})$. Hence, to compare the convergence rate of Muon and GD, we need to compare $LD_{\rm F}^{2}$ with $L_{*}D_{{\rm op}}^{2}$. Similarly, if $f$ satisfies 4.11 with $\Lambda$ such that $\|\Lambda\|_{{\rm op}}\approx\|\Lambda\|_{*}\ll r\|\Lambda\|_{{\rm op}}$, then $L_{*}D_{{\rm op}}^{2}\leq\|\Lambda\|_{*}D_{{\rm op}}^{2}\ll r\|\Lambda\|_{{\rm op}}D_{{\rm op}}^{2}$. In the experiments, we usually find that $\|W_{t}-W^{*}\|_{{\rm op}}$ decreases as $t$ increases (Figure 1(b)). Thus, we assume $D_{{\rm op}}\approx\|W_{0}-W^{*}\|_{{\rm op}}$. Despite the success of LoRA (Hu et al., 2022) in fine-tuning foundation models indicating a relatively low rank of $W_{0}-W^{*}$ is enough for fine-tuning, many empirical evidences (Lialin et al., 2024; Jiang et al., 2024; Huang et al., 2025) demonstrate that a relatively high rank of $W_{0}-W^{*}$ is needed for pretraining and some complex fine-tuning tasks for large foundation models, and a low rank of $W_{0}-W^{*}$ can lead to degraded performance in those scenarios. Thus, we assume $W_{0}-W^{*}$ is relatively “high rank”. Then, we can expect $r\|W_{0}-W^{*}\|^{2}_{{\rm op}}\approx\|W_{0}-W^{*}\|^{2}_{\rm F}$ and $L_{*}D_{{\rm op}}^{2}\ll r\|\Lambda\|_{{\rm op}}D_{{\rm op}}^{2}\approx LD_{\rm F}^{2}$, which means Muon can have a better performance than GD.
To further evaluate this argument, we conduct experiments comparing Muon and GD on a quadratic function $f(W)=\text{tr}\left((W-W^{*})^{\top}Q(W-W^{*})\right)$, where $Q$ is a positive definite matrix with a poor condition number and is relatively “low rank”, i.e. $\|Q\|_{*}\approx\|Q\|_{{\rm op}}$. Note that the Hessian of $f$ is $I\otimes Q$, which satisfies 4.11. We randomly choose $W^{*}$, and set $W_{0}=0$. We report the results in Figure 1. We can see that Muon outperforms GD and most samples of ratio $\frac{D_{\rm F}^{2}L}{D_{{\rm op}}^{2}L_{*}}$ are larger than 3, which validates the theoretical findings.
(a) Function Value
(b) $\|W_{t}-W^{*}\|_{{\rm op}}$
(c) $D_{\rm F}^{2}L/(D_{{\rm op}}^{2}L_{*})$
Figure 1: Experiments on a quadratic function $f(W)=\text{tr}\left((W-W^{*})^{\top}Q(W-W^{*})\right)$. Detailed settings can be found in Appendix E.
Next, we consider a more realistic example to illustrate how the low-rank structure of the Hessian arises and how Muon can benefit from it. Specifically, we consider a multi-class classification problem with dataset $\{(x_{i},y_{i})\}^{B}_{i=1}$, where $x_{i}\in\mathbb{R}^{d}$ is the feature data, $y_{i}\in\mathbb{R}^{c}$ is the corresponding $c$-dimensional one-hot vector of the label, $c$ is the number of classes and $B$ is the number of samples. We first consider a linear model $g(W;x)=Wx\in\mathbb{R}^{c}$, where $W\in\mathbb{R}^{c\times d}$ is the parameter, and its mean-square (MSE) loss. Denote $X=(x_{1},\dots,x_{B})\in\mathbb{R}^{d\times B}$, $Y=(y_{1},\dots,y_{B})\in\mathbb{R}^{c\times B}$. The MSE loss can be defined as follows:
$f(W)=\frac{1}{2B}\|WX-Y\|_{\rm F}^{2}.$ (3)
We can calculate the Hessian of this loss as
$\nabla^{2}f_{v}(\overline{\mathrm{vec}}(W))=I_{c}\otimes XX^{\top}.$ (4)
We note that this Hessian is block-diagonal with $c$ blocks and satisfy 4.11 with $\Lambda=XX^{\top}$. We can calculate $L=\|XX^{\top}\|_{\rm op}=\|X\|_{\rm op}^{2}$ and $L_{*}=\|XX^{\top}\|_{*}=\|X\|_{\rm F}^{2}$.
(a) Spectra of $X_{\text{MNIST}}$
(b) Spectra of $X_{\text{Gaussian}}$
(c) $X_{\text{MNIST}},\,W\in\mathbb{R}^{10\times 784}$. Ratio: $D_{\rm F}^{2}L/(D_{{\rm op}}^{2}L_{*})=1.57$
(d) $X_{\text{MNIST}},\,W\in\mathbb{R}^{100\times 784}$. Ratio: $D_{\rm F}^{2}L/(D_{{\rm op}}^{2}L_{*})=3.52$
(e) $X_{\text{Gaussian}},\,W\in\mathbb{R}^{10\times 784}$. Ratio: $D_{\rm F}^{2}L/(D_{{\rm op}}^{2}L_{*})=0.0122$
(f) $X_{\text{Gaussian}},\,W\in\mathbb{R}^{100\times 784}$. Ratio: $D_{\rm F}^{2}L/(D_{{\rm op}}^{2}L_{*})=0.0232$
Figure 2: (a, b): Spectra of $X_{\text{MNIST}}$ and $X_{\text{Gaussian}}$. (c, d, e, f): Optimizing (3) with different $X$ and $Y$.
Actually, we find that the feature matrix $X$ in practical machine learning problem is typically very “low rank”, or more precisely, their singular values are highly concentrated (See Figure 2(a) and additional examples in Appendix D) such that $\|X\|_{\rm op}\approx\|X\|_{\rm F}$. For example, we randomly select 1000 samples from training datasets of MNIST (LeCun et al., 1998) and construct $X_{\text{MNIST}}\in\mathbb{R}^{784\times 1000}$. Our calculation shows that $\|X_{\text{MNIST}}\|^{2}_{\rm F}/\|X_{\text{MNIST}}\|^{2}_{\rm op}=1.41$, which means $L_{*}\approx L$, the Hessian of this loss is extremely low-rank. Note that this phenomenon is not limited to MNIST; similar low-rank behavior can also be observed in other real-world datasets (e.g., CIFAR-10 and text data), see Appendix D.
According to the theory, Muon can benefit from this low-rank structure and potentially converge faster than GD, thus, to verify this, we conduct the following experiments. We first use $X_{\text{MNIST}}$ and their corresponding 1000 labels $y$ from the training dataset as $Y\in\mathbb{R}^{10\times 1000}$ to optimize $W\in\mathbb{R}^{10\times 784}$ under the MSE loss in (3). We compare GD (with/without Nesterov momentum), Adam, and Muon, tuning all hyperparameters for each method. Results in Figure 2(c) show that Muon outperforms GD and Adam. Then, we compute the Frobenius norm and operator norm of $W_{0}-W^{*}$ (where $W^{*}$ is selected as the final converged point), denoted as $D_{F}$ and $D_{\text{op}}$, respectively, and calculate the ratio $\frac{D_{\rm F}^{2}L}{D_{{\rm op}}^{2}L_{*}}$. This ratio equals 1.57, exceeding 1 and thus supporting Muon’s advantage, though it is not large, explaining the modest empirical gain in Figure 2(c). One reason is that the intrinsic dimension of $W\in\mathbb{R}^{10\times 784}$ limits its rank to be relatively low, which in turn leads to a relatively small value of $D_{\mathrm{F}}^{2}/D_{\mathrm{op}}^{2}$. To amplify this effect, we replace $Y$ with 1000 random one-hot vectors over 100 classes, yielding $W\in\mathbb{R}^{100\times 784}$. We then conduct the same experiment comparing GD, Adam, and Muon. The results are shown in Figure 2(d). We can observe that the ratio is now larger, reaching $3.52$. Correspondingly, Muon more significantly outperforms GD and Adam in the figure, which is consistent with our theoretical predictions.
While, if we construct $X_{\text{Gaussian}}\in\mathbb{R}^{784\times 1000}$ as a Gaussian random matrix, then we can find that the singular values of this random matrix are more evenly distributed (See Figure 2(b)) and $\|X_{\text{Gaussian}}\|^{2}_{\rm F}/\|X_{\text{Gaussian}}\|^{2}_{\rm op}=224.06$, which means its $L_{*}$ is significantly larger than $L$ and indicating Muon will outperform worser than GD. Our experiments (Figure 2(e, f)) validated our theoretical predictions. Notably, when we increase the dimension of $y$ from 10 to 100, the corresponding ratio rises from 0.0122 to 0.0232, and the performance gap between Muon and GD (Adam) becomes smaller. As shown in Figure 2(c, d, e, f), Muon performs better relative to GD when this ratio is larger, which is consistent with our theoretical analysis. Meanwhile, our experiments also suggest that the low-rank property of the Hessian in practical machine learning problems may partially stem from the fact that real-world data are not Gaussian random, but instead exhibit some inherent low-rank structure.
4.3 Without uniform Lipschitz smoothness assumptions
@@
Though the previous subsections offer some insight into how Muon can outperform GD, 4.11 appears overly restrictive. To further investigate how the structure and dynamics of the Hessian matrix affect the convergence of Muon and GD in a more fine-grained manner, we consider the following assumption and analysis.
@@
We assume the norm of $\nabla^{3}f$ is bounded, i.e., we assume for any $W\in\mathbb{R}^{m\times n}$, $\|\nabla^{3}f_{v}(w)\|_{{\rm op}}\leq s$. Here $f_{v}(w)=f(W)$ and $w=\overline{\mathrm{vec}}(W)\in\mathbb{R}^{mn}$.
@@
Under 4.13, if we apply Algorithm 2 with $\eta_{t}=\eta$, then
@@
where $J=\frac{1}{T}\sum_{t=0}^{T-1}J_{t}$, $J_{t}=\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})^{\top}H_{t}\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})$, $H_{t}=\nabla^{2}f_{v}(\overline{\mathrm{vec}}(W_{t}))$, and $f_{v}(\overline{\mathrm{vec}}(W_{t}))=f(W_{t})$.
If J>0J>0111We find that usually $J_{t}>0$ in experiments. Thus, we mainly discuss the situations when $J>0$ ($\tilde{J}>0$) in the main paper. When $J\leq 0$ ($\tilde{J}\leq 0$), there is a better convergence rate regarding $T$; see discussions in Appendix B.3 and C.5., and $\eta=O\left(\sqrt{\frac{\Delta}{JT}}\right)$, then
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla f(W_{t})\|_{*}\leq\left(\sqrt{\frac{J\Delta}{T}}+\frac{sr^{3/2}\Delta}{JT}\right).$
Thus, it is possible that Muon can find an $\epsilon$-nuclear norm stationary point of $f$ with a complexity of $O(J\Delta\epsilon^{-2})$.
Similarly, for the star convex case, we have
Theorem 4.15.
Under Assumptions 4.3, 4.13 and 4.4, if we apply Algorithm 2 with $\eta_{t}=\eta$, we have
$f(W_{T})-f^{*}$ $\leq\left(1-\frac{\eta}{D_{{\rm op}}}\right)^{T}\Delta+\frac{\eta^{2}\tilde{J}T}{2}+\frac{s\eta^{2}D_{\rm op}r^{3/2}}{6}.$
where $\tilde{J}=\frac{1}{T}\sum_{t=0}^{T-1}\left(1-\frac{\eta}{D_{{\rm op}}}\right)^{T-1-t}J_{t}$, $J_{t}=\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})^{\top}H_{t}\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})$, $H_{t}=\nabla^{2}f_{v}(\overline{\mathrm{vec}}(W_{t}))$, and $f_{v}(\overline{\mathrm{vec}}(W_{t}))=f(W_{t})$.
Note that $\tilde{J}$ can be viewed as a weighted average of $J_{t}$ and $\tilde{J}\leq\frac{1}{T}\sum_{t=0}^{T-1}|J_{t}|$. If $\tilde{J}>0$1 and $\eta=\min\left\{\frac{D_{{\rm op}}}{T}\log\left(\frac{T\Delta}{D_{{\rm op}}^{2}\tilde{J}}\right),D_{\rm op}\right\}$, then
$f(W_{T})-f^{*}\leq$ $\frac{D_{{\rm op}}^{2}\tilde{J}}{T}+\frac{D_{{\rm op}}^{2}\tilde{J}}{2T}\left[\log\left(\frac{T\Delta}{D_{{\rm op}}^{2}\tilde{J}}\right)\right]^{2}+\frac{sr^{3/2}D_{{\rm op}}^{3}}{6T^{2}}\left[\log\left(\frac{T\Delta}{D_{{\rm op}}^{2}\tilde{J}}\right)\right]^{2}$
$\leq$ $\tilde{O}\left(\frac{D_{{\rm op}}^{2}\tilde{J}}{T}\right).$
Thus, it is possible that Muon can reach the precision $f(W_{T})-f^{*}\leq\epsilon$ with a complexity of $\tilde{O}(\tilde{J}D_{{\rm op}}^{2}\epsilon^{-1})$.
4.3.1 Comparison with GD
Therefore, the key to comparing Muon and GD lies in comparing the relationship between $J$ (or $\tilde{J}$) and $L$. Note that $J$ is the average of $J_{t}$, while $\tilde{J}$ is bounded above by the average of $|J_{t}|$. Hence, we may first analyze $J_{t}$. Note that
$J_{t}$ $=\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})^{\top}H_{t}\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})$
$=\overline{\mathrm{vec}}(I_{r_{t}})^{\top}(U_{t}^{\top}\otimes V_{t}^{\top})H_{t}(U_{t}\otimes V_{t})\overline{\mathrm{vec}}(I_{r_{t}})$